The sum of ages of man and wife is 6 times the ages of the children.Two years ago the sum was 10 times the sum of ages of their children at that time.After six years sum of their children will be 3 times the sum of ages of their children.How many children do the have?

Grade:10

1 Answers

aritra nandy
13 years ago

let mans age be a,womans be b,children be c,

a+b=6c

a-2+b-2=10c

a+b-4=10c

a+b+12=3c

now use crammers rule